OpenStudy (anonymous):

need help on logarithmic equation 12^(x-2)=20

OpenStudy (unklerhaukus):

12^(x-2) = 20 take the log to base 12 on both sides log_12[12^(x-2)] = log_12[20] (x-2)log_12[12] = log_12[20]

OpenStudy (anonymous):

ok i understand how to set it up like that but what do i do from there

OpenStudy (unklerhaukus):

x-2= log_12[20]

OpenStudy (anonymous):

so i would enter it in the calculator how? sorry im not too good at math

OpenStudy (unklerhaukus):

the log term dosen't simplify just get x by itself on one side

OpenStudy (anonymous):

so it would be x=2??

OpenStudy (unklerhaukus):

where did the log go?

OpenStudy (unklerhaukus):

x-2= log_12[20] add two to both sides

OpenStudy (unklerhaukus):

x =

OpenStudy (anonymous):

20

OpenStudy (unklerhaukus):

? whgat

OpenStudy (anonymous):

??? whats the answer lol it would make much better sense

OpenStudy (unklerhaukus):

if you try that solution in the original equation, the left hand side would become 12^(20-2) 12^18 but this is much more than the right hand side 20

OpenStudy (anonymous):

30? lol

OpenStudy (unklerhaukus):

try again

OpenStudy (unklerhaukus):

the correct answer has a log in it

OpenStudy (anonymous):

log 12

OpenStudy (unklerhaukus):

guessing isn't an effective method for maths

OpenStudy (anonymous):

no clue?

OpenStudy (unklerhaukus):

if there is something i have said that you do not understand please ask about it

OpenStudy (anonymous):

i dont undertstand this --> x-2= log_12[20] add two to both sides

OpenStudy (unklerhaukus):

x-2= log_12[20] adding two to both sides x-2+2= log_12[20]+2 the left hand side simplifies